[2] Your answer is partially correct. Try again. An electrically neutral model airplane is flying in a horizontal circle on a 3.5-m guideline, which is nearly parallel to the ground. The line breaks when the kinetic energy of the plane is 50 1. Reconsider the same situation, except that now there is a point charge of +q on the plane and a point charge of -q at the other end of the quideline. In this case, the line breaks when the kinetic energy of the plane is 52.0 J. Find the magnitude of the charges 2.8e-5

Two spherical objects are separated by a distance of 2.01 x 10 m. The objects are initially electrically neutral and are very small compared to the distance between them. Each object acquires the same negative charge due to the addit ono electrons. As a result, each object experiences an electrostatic rce that has a magnitude of 2 S 10 N Ho many eet ons did it take to produce the charge on one of the objects? NumberTT6e-38 No units V exact number, no tolerance
Answer #1

Solution :- Given data:-

kinetic energy is mv2/2 =52 J

           k is constant = 9*10^9 Nm2/C2
            q1and q2 are the charge of the particles
circular radius = 3.5 m
         since q1= q2 =q

    when a particle is moving in a horizontal circle its centrifugal force is equivalent to electrostatic force between particle and the particle at the center
       so q192
         by multiplying with 1/2 on both sides and cancelling the r in the denominator
                  1 29142

52. = [(9×10^9)×(q^2)]÷[2×3.5]

q = 2.011× 10^-4 Columbus.